# AADNet: An End-to-End Deep Learning Model for Auditory Attention Decoding

### v1

@@ Abstract
Auditory attention decoding (AAD) is the process of identifying the attended speech in a multi-talker environment using brain signals, typically recorded through electroencephalography (EEG). Over the past decade, AAD has undergone continuous development, driven by its promising application in neuro-steered hearing devices. Most AAD algorithms are relying on the increase in neural entrainment to the envelope of attended speech, as compared to unattended speech, typically using a two-step approach. First, the algorithm predicts representations of the attended speech signal envelopes; second, it identifies the attended speech by finding the highest correlation between the predictions and the representations of the actual speech signals. In this study, we proposed a novel end-to-end neural network architecture, named AADNet, which combines these two stages into a direct approach to address the AAD problem. We compare the proposed network against the traditional approaches, including linear stimulus reconstruction, canonical correlation analysis, and an alternative non-linear stimulus reconstruction using two different datasets. AADNet shows a significant performance improvement for both subject-specific and subject-independent models. Notably, the average subject-independent classification accuracies from 56.1% to 82.7% with analysis window lengths ranging from 1 to 40 seconds, respectively, show a significantly improved ability to generalize to data from unseen subjects. These results highlight the potential of deep learning models for advancing AAD, with promising implications for future hearing aids, assistive devices, and clinical assessments.
@@ 1 Introduction
In a noisy environment, the human brain demonstrates a remarkable ability to segregate sound streams, allowing individuals to focus on the sound of interest while disregarding the others. For individuals with hearing impairment, this ability is often significantly deteriorated [1]. Assistive hearing devices equipped with noise suppression and speech enhancement algorithms can partially compensate for this deficit. However, these hearing devices tend to perform poorly in a multi-talker environment due to a lack of information about the target talker. Identifying the attended talker, i.e., auditory attention decoding (AAD), has garnered significant attention from researchers over the past decades due to potential applications in future neuro-steered assistive hearing devices.
@@
Previous studies have demonstrated evidence of neural encoding of different speech features through various recording methods, such as the speech envelope via EEG signals [2], spectrogram via electrocorticography (ECoG) [3], and magnetoencephalography (MEG) [4, 5, 6]. As EEG is minimally invasive and can be integrated into portable devices, following the study of Aiken et al.[2], the majority of AAD methods are based on the envelope following response extracted from EEG signals, and these methods have successfully demonstrated AAD across various experimental paradigms [7, 8, 9, 10, 11, 12, 13, 14, 15]. This approach has become a dominant and well-established method for addressing the AAD problem. While the aforementioned methods rely on acoustical features, another explored linguistic and lexical features, such as the onset and surprisal of words and phonemes, to decode neural responses to speech [16]. These approaches assume that words and phonemes are annotated and that language-specific dictionaries with word and phoneme probabilities are accessible. Additionally, Raghavan et al.[17] proposed a system that identifies auditory events using the local maxima in the envelope rate of change, and utilized masking-specific event-related potential classifiers to determine the attended sound source, suggesting a new approach to AAD.
@@
The most common approach used in envelope-based AAD algorithms is backward modeling in which a decoder is trained to reconstruct the attended speech envelope. During the training procedure, the Pearson correlation or the mean square error between the reconstructed and actual envelopes is used as the objective function to optimize the model’s parameters. The reconstructed envelope is then correlated with the actual envelopes and used as input to an additional classifier to determine the attended speech stream. Many studies have used linear decoders across various recording paradigms [7, 8, 9, 10], while others have employed non-linear models with different neural network structures, such as fully connected neural networks (FCNN) [13, 15], convolutional neural networks (CNN) [15], long-short term memory (LSTM) [18], and CNN-LSTM [19], achieving promising results. Generally, the linear models seem to be consistent and well-established due to their simplicity (a low number of parameters) and have been applied across various datasets. The non-linear counterparts typically have large numbers of parameters and have been reported to outperform the linear models [15, 18] due to their capacity to model the nonlinearity of speech processing in the auditory system. However, most of these non-linear methods are designed and validated on specific paradigms from the original studies, and they may not generalize well to different datasets or recording paradigms. An alternative approach for backward modeling is forward modeling, which predicts the EEG response from the speech envelopes. The predicted EEG is then used to compare with the measured EEG to determine the attended speech. This approach has been reported to underperform compared to the backward models [20]. This result is understandable, given the challenge of predicting neural responses, as they result from a combination of various neural activities triggered by multiple internal and external stimuli, with only the audio being known. Other studies have attempted a combined forward-backward approach using the Canonical Component Analysis (CCA) algorithm [12, 21] to convert both EEG and envelope signals into maximized correlated latent vectors, followed by a classifier to determine the attended speech. This approach has been demonstrated to outperform the other linear decoder methods [22].
@@
Common to the approaches mentioned above is that they involve two separate stages: training models to convert the EEG signals and/or the audio envelopes into latent vectors, calculating the correlation scores, and using a classifier to decode the attention. Recently, Ciccarelli et al.[14] have proposed a direct approach using a CNN model that does not explicitly reconstruct the envelope. This new approach has been shown to outperform the previous linear and non-linear envelope reconstruction approaches. However, like the majority of other AAD studies, this approach has not been tested for subject-independent models which is an important yet underdeveloped topic, as pointed out by other authors [23].
@@
This study adopts the direct approach by proposing AADNet, an end-to-end deep learning (DL) model to address the AAD problem. The model utilizes EEG signals and the audio envelopes of two speakers, to directly determine the attended speaker without reconstructing the attended envelopes. To evaluate the performance of the proposed model, we conduct a comparison with other state-of-the-art AAD methods: linear stimulus reconstruction (LSR) [8], CCA [12], and non-linear stimulus reconstruction (NSR) [15] on two datasets for both subject-specific (SS) and subject-independent (SI) models.
@@ 2 Envelope-based AAD algorithms
As previously mentioned, AAD algorithms predominantly focus on the correlation between the recorded EEG signals and the envelope of the attended stimulus. According to Geirnaert et al.[22], the most robust AAD methods are the LSR and the CCA model that combines the forward and backward approaches. Therefore, in this study, we implemented these methods as least-squares-based baselines to compare with the proposed method.
The concept of decoding attention based on a stimulus reconstruction approach is depicted in Figure 1. First, multi-channel EEG signals and audio signals from each stream are pre-processed (see Section 3.2). Subsequently, the actual envelopes of the audio signals in each audio stream are extracted, assuming the demixed speech streams are available. In the training phase, the envelope of the attended stream and the EEG signals are used to train a decoder. In the inference phase, the EEG signals are used to reconstruct the envelope of the attended stream. The attended stream is determined as the one whose envelope is most highly correlated with the reconstructed envelope. The decoder can be constructed as a linear or non-linear model that maps the EEG data to the attended or unattended audio envelope. However, as found by O’Sullivan et al.[8], the attended decoder obtains higher decoding accuracy than the unattended one does. Therefore, in this study, we use the attended decoder.
@@ 2.1.1 Linear stimulus reconstruction
where $\mathbf{g}\in\mathbb{R}^{NL\times 1}$ is the decoder, collecting all decoder coefficients for all time lags and channels. Assuming that there are $T$ envelope samples available, $\mathbf{s}\in\mathbb{R}^{1\times T}$ is the envelope vector, $\mathbf{X}\in\mathbb{R}^{NL\times T}$ is the EEG matrix with each column vector contains all the EEG signals of all time lags and channels, $\mathbf{I}$ is the identity matrix, and $\lambda$ is the regularization parameter, which is estimated from a cross-validation approach from a set of values ranging from $10^{-2}$ to $10^{10}$ with a logarithmic step. The time lag $\tau$ covers a temporal EEG context from 0 up to 250 ms post-stimulus, as it has been found to have the best decoding accuracy [8].
@@ 2.1.2 CCA
CCA is a multivariate statistical technique used to analyze the relationship between two sets of variables [24]. The CCA method has been successfully applied to the AAD problem and has achieved promising results [12, 21]. In AAD, the goal of CCA is to find an optimal spatiotemporal linear transform (decoder) $\mathbf{w}_{x}\in\mathbb{R}^{NL\times 1}$ on EEG signals $\mathbf{X}\in\mathbb{R}^{NL\times{T}}$ and a temporal linear transform (encoder) $\mathbf{w}_{s}\in\mathbb{R}^{L_{s}\times 1}$ on audio envelopes $\mathbf{S}\in\mathbb{R}^{L_{s}\times T}$ to maximize the correlation between the two latent vectors. $L$ and $L_{s}$ correspond to the decoder length and encoder length, respectively. The CCA method can be described as the following optimization equation:
$\mathbf{\hat{w}}_{x},\mathbf{\hat{w}}_{s}$ $=\operatorname*{argmax\;}_{\mathbf{w}_{x},\mathbf{w}_{s}}{\frac{\mathbf{w}_{x}^{\textsc{t}}\mathbf{R_{xs}}\mathbf{w}_{s}}{\sqrt{\mathbf{w}_{x}^{\textsc{t}}\mathbf{R_{xx}}\mathbf{w}_{x}}\sqrt{\mathbf{w}_{s}^{\textsc{t}}\mathbf{R_{ss}}\mathbf{w}_{s}}}}.$ (3)
@@ 2.2.1 Non-linear stimulus reconstruction
Another approach is reconstructing the stimulus envelope using a non-linear model (NSR). In this study, we implemented the CNN-based network proposed by Thorton et al.[15] as a baseline method for the NSR approach. The network was inspired by the EEGNet architecture developed by Lawhern et al.[25], which comprises two main convolutional blocks, one fully connected (FC) classification layer and employs the exponential linear unit (ELU) as a nonlinear activation function, as well as batch normalization (BN) and average pooling. Details of the network can be found in the original study [15]. Another notable study using the NSR approach is by De Taillez et al.[13]. The authors developed a feed-forward neural network comprising a single hidden layer and an output layer with other DL features such as ‘tanh’ activation functions, dropout (DO) [26], and BN. Despite our best efforts to implement and validate this method, its performance was significantly lower compared to the other methods, and therefore, the results are not included here. A similar observation was reported in [22].
@@ 2.2.2 Proposed direct AAD method
Figure 3: The architecture of AADNet. T = length of input data, C = number of input EEG data, BN = batch normalization, DO = dropout, FC = fully connected, and h = output size of the FC1 layer.
@@ 3.1 Dataset
In this work, we use two datasets recorded using a competing-talker setup, which has been used in previous AAD-related studies to address the AAD problem.
@@ 3.1.2 Dataset II - DTU
Table 2 summarizes the details of the two datasets.
@@
DTU 18 60 50 64 Insert earphones, simulated speaker direction of ±${\pm}60^{\circ}$ through a head-related transfer function, 2.4 m distance Anechoic, mildly, and highly reverberant
@@ 3.2 Data pre-processing
In this section, we present the processing steps that were applied on the two presented datasets for the LSR, CCA, and the proposed AADNet models. For the NSR model, we followed the processing steps proposed in the original study [15].
@@ 3.2.1 EEG
The EEG data were first band-pass filtered using a zero-phase FIR filter with cutoffs at 0.5 Hz and 32 Hz to eliminate slow drift and irrelevant high frequencies. This frequency band was also applied in various studies [13, 14, 22] and is reported to be relevant for envelope-based AAD studies. The data for each channel were then downsampled to 64 Hz, re-referenced to the average of all channels (32 channels for the EventAAD dataset and 64 channels for the DTU dataset), and zero-centered. For the NSR model, the EEG data were band-pass filtered from 0.25 to 36 Hz using a Hamming window, FIR filter, resampled to 125 Hz, and standardized to have zero mean and unit variance. The pre-processing pipeline was implemented using the Python MNE package version 1.2.0 [31] and SciPy package version 1.10.1 [32].

### v2

@@ Abstract
Auditory attention decoding (AAD) is the process of identifying the attended speech in a multi-talker environment using brain signals, typically recorded through electroencephalography (EEG). Over the past decade, AAD has undergone continuous development, driven by its promising application in neuro-steered hearing devices. Most AAD algorithms are relying on the increase in neural entrainment to the envelope of attended speech, as compared to unattended speech, typically using a two-step approach. First, the algorithm predicts representations of the attended speech signal envelopes; second, it identifies the attended speech by finding the highest correlation between the predictions and the representations of the actual speech signals. In this study, we proposed a novel end-to-end neural network architecture, named AADNet, which combines these two stages into a direct approach to address the AAD problem. We compare the proposed network against traditional stimulus decoding-based approaches, including linear stimulus reconstruction, canonical correlation analysis, and an alternative non-linear stimulus reconstruction using three different datasets. AADNet shows a significant performance improvement for both subject-specific and subject-independent models. Notably, the average subject-independent classification accuracies for different analysis window lengths range from 56.3% (1 s) to 78.1% (20 s), 57.5% (1 s) to 89.4% (40 s), and 56.0% (1 s) to 82.6% (40 s) for three validated datasets, respectively, showing a significantly improved ability to generalize to data from unseen subjects. These results highlight the potential of deep learning models for advancing AAD, with promising implications for future hearing aids, assistive devices, and clinical assessments.
@@ 1 Introduction
The human brain demonstrates a remarkable ability to segregate sound streams, allowing individuals to focus on the sound of interest while disregarding the others. For individuals with hearing impairment, this ability is often significantly deteriorated [1]. Assistive hearing devices equipped with noise suppression and speech enhancement algorithms can partially compensate for this deficit. However, these hearing devices tend to perform poorly in a multi-talker environment due to a lack of information about the target talker. Therefore, over the past decade, various research has been conducted to address this challenge by identifying the attended talker from brain signals, i.e., auditory attention decoding (AAD). This can potentially lead to the development of so-called neuro-steered hearing devices, in which the attended speech is enhanced by suppressing irrelevant sound sources. While this paper focuses specifically on the AAD challenge itself, alternative approaches in neuro-steered hearing devices address this issue by simultaneously performing speech separation and enhancement informed by brain signals, known as brain-informed speech separation [2, 3].
@@
Previous studies have demonstrated neural encoding of various speech features using different recording techniques, including the speech envelope via EEG [4], spectrogram via electrocorticography (ECoG) [5], and via magnetoencephalography (MEG) [6, 7, 8]. As EEG is minimally invasive and can be integrated into portable devices, following the study of Aiken et al.[4], the majority of AAD methods are based on the envelope following response extracted from EEG signals, and these methods have successfully demonstrated AAD across various experimental paradigms [9, 10, 11, 12, 13, 14, 15, 16, 17]. This approach has become a dominant and well-established method for addressing the AAD problem. While the aforementioned methods rely on acoustical features, another approach has investigated linguistic and lexical features, such as the onset and surprisal of words and phonemes, to decode neural responses to speech [18]. These approaches assume that words and phonemes are annotated and that language-specific dictionaries with word and phoneme probabilities are accessible. Additionally, Raghavan et al.[19] proposed a system that identifies auditory events using the local maxima in the envelope rate of change, and utilized masking-specific event-related potential classifiers to determine the attended sound source, suggesting a new approach to AAD.
@@
The most common approach used in envelope-based AAD algorithms is backward modeling in which a decoder is trained to reconstruct the attended speech envelope. During the training procedure, the Pearson correlation or the mean square error between the reconstructed and actual envelopes is used as the objective function to optimize the model’s parameters. The reconstructed envelope is then correlated with the actual envelopes and used as input to an additional classifier to determine the attended speech stream. Many studies have used linear decoders across various recording paradigms [9, 10, 11, 12], while others have employed non-linear models with different neural network structures, such as fully connected neural networks (FCNN) [15, 17], convolutional neural networks (CNN) [17], long-short term memory (LSTM) [20], and CNN-LSTM [21], achieving promising results. Generally, the linear models seem to be consistent and well-established due to their simplicity (a low number of parameters) and have been applied across various datasets. The non-linear counterparts typically have large numbers of parameters and have been reported to outperform the linear models [17, 20] due to their capacity to model the nonlinearity of speech processing in the auditory system. However, most of these non-linear methods are designed and validated on specific paradigms from the original studies, and they may not generalize well to other datasets or recording paradigms.
An alternative approach to backward modeling is forward modeling, which predicts the EEG response from the speech envelopes. The predicted EEG is then used to compare with the measured EEG to determine the attended speech. This approach has been reported to underperform compared to the backward models [22]. This result appears reasonable, given that neural responses arise from a large number of underlying neural sources, of which only a small subset are directly or indirectly related to the audio. Other studies have attempted a combined forward-backward approach using the Canonical Component Analysis (CCA) algorithm [14, 23] to convert both EEG and envelope signals into maximized correlated latent vectors, followed by a classifier to determine the attended speech. This approach has been demonstrated to outperform the other linear decoder methods [24].
@@
Common to the approaches mentioned above is that they involve two separate stages: training models to convert the EEG signals and/or the audio envelopes into latent vectors, calculating the correlation scores, and using a classifier to decode the attention. Recently, Ciccarelli et al.[16] have proposed a direct approach using a CNN model that does not explicitly reconstruct the envelope. This new approach has been shown to outperform the previous linear and non-linear envelope reconstruction approaches. However, like the majority of other AAD studies, this approach has not been tested for subject-independent models, which is an important yet underdeveloped topic, as pointed out by other authors [25].
@@
The motivation for this study is to improve AAD performance, specifically, its generalization capability to unseen subjects, taking a step towards integrating AAD algorithms into real-life applications. To achieve this, we adopt the direct approach by proposing AADNet, an end-to-end deep learning (DL) model to address the AAD problem. The model utilizes EEG signals and the audio envelopes of two speakers to directly determine the attended speaker without reconstructing the attended envelopes. To evaluate the performance of the proposed model, we conduct a comparison with other state-of-the-art AAD methods: linear stimulus reconstruction (LSR) [10], CCA [14], and non-linear stimulus reconstruction (NSR) [17] on three datasets for both subject-specific (SS) and subject-independent (SI) models. The results demonstrate that the proposed model outperforms the other methods, particularly in the subject-independent setting. The code is publicly available at https://github.com/babibo180918/AADNet.
@@ 2 Envelope-based AAD algorithms
As previously mentioned, current AAD algorithms predominantly focus on the correlation between the recorded EEG signals and the envelope of the attended stimulus. According to Geirnaert et al.[24], the most robust AAD methods are the LSR and the CCA model that combines the forward and backward approaches. Therefore, in this study, we implemented these methods as least-squares-based baselines to compare with the proposed method. Apart from that, with breakthroughs across multiple disciplines, deep learning (DL) methods have been introduced in various studies to advance the field [15, 16, 26, 21, 17, 20, 18, 27]. However, these methods address different tasks (single-speaker match-mismatch, AAD speaker identification (SpkI), and locus of attention (LoA)), utilize different recording modalities (EEG, MEG), and employ various speech features (linguistics, speech envelope, spectrogram). A direct comparison of all these methods is beyond the scope of this study. In this study, we include DL-based studies that meet the following criteria:
  • Recent DL-based model, where the original paper proposing the model did not utilize both the two publicly available datasets DTU and KUL (see Section 3.1.2 and Section 3.1.3).
  • Methods designed for the AAD speaker identification task using envelope features.
  • Methods with publicly available source code.
Typically, other studies use specific datasets to develop and tune their proposed models. This study follows a similar approach by using the EventAAD dataset (see Section 3.1.1) to develop our model, and only using the DTU and KUL datasets for validating the model. Therefore, the first criterion ensures that the hyperparameters of the included models were not tailored to the datasets used in the model comparison. Based on these criteria, the study by Thornton et al.[17] is considered the state-of-the-art for NSR methods.
The concept of decoding attention based on a stimulus reconstruction approach is depicted in Figure 1. First, multi-channel EEG signals and audio signals from each stream are pre-processed (see Section 3.2). Subsequently, the actual envelopes of the audio signals in each audio stream are extracted, assuming the demixed speech streams are available. In the training phase, the envelope of the attended stream and the EEG signals are used to train a decoder. In the inference phase, the EEG signals are used to reconstruct the envelope of the attended stream. The attended stream is determined as the one whose envelope has the highest correlation with the reconstructed envelope. The decoder can be constructed as a linear or non-linear model that maps the EEG data to the attended or unattended audio envelope. However, as found by O’Sullivan et al.[10], the attended decoder obtains higher decoding accuracy than the unattended one does. Therefore, in this study, we use the attended decoder.
@@ 2.1.1 Linear stimulus reconstruction
where $\mathbf{g}\in\mathbb{R}^{NL\times 1}$ is the decoder, collecting all decoder coefficients for all time lags and channels. Assuming that there are $T$ envelope samples available, $\mathbf{s}\in\mathbb{R}^{T\times 1}$ is the envelope vector, $\mathbf{X}\in\mathbb{R}^{T\times NL}$ is the EEG matrix with each row vector contains all the EEG signals across all time lags and channels, $\mathbf{I}$ is the identity matrix, and $\lambda$ is the regularization parameter, which is estimated from a cross-validation approach from a set of values ranging from $10^{-2}$ to $10^{10}$ with a logarithmic step. The time lag $\tau$ covers a temporal EEG context from 0 up to 250 ms post-stimulus, as it has been found to have the best decoding accuracy [10].
@@ 2.1.2 CCA
CCA is a multivariate statistical technique used to analyze the relationship between two sets of variables [28]. The CCA method has been successfully applied to the AAD problem and has achieved promising results [14, 23]. In AAD, the goal of CCA is to find an optimal spatiotemporal linear transform (decoder) $\mathbf{w}_{x}\in\mathbb{R}^{NL\times 1}$ on EEG signals $\mathbf{X}\in\mathbb{R}^{T\times{NL}}$ and a temporal linear transform (encoder) $\mathbf{w}_{s}\in\mathbb{R}^{L_{s}\times 1}$ on audio envelopes $\mathbf{S}\in\mathbb{R}^{T\times L_{s}}$ to maximize the correlation between the two latent vectors. $L$ and $L_{s}$ correspond to the decoder length and encoder length, respectively. The CCA method can be described as the following optimization equation:
$\mathbf{\hat{w}}_{x},\mathbf{\hat{w}}_{s}$ $=\operatorname*{argmax\;}_{\mathbf{w}_{x},\mathbf{w}_{s}}{\frac{\mathbf{w}_{x}^{\textsc{t}}\mathbf{R_{xs}}\mathbf{w}_{s}}{\sqrt{\mathbf{w}_{x}^{\textsc{t}}\mathbf{R_{xx}}\mathbf{w}_{x}}\sqrt{\mathbf{w}_{s}^{\textsc{t}}\mathbf{R_{ss}}\mathbf{w}_{s}}}},$ (3)
where $\mathbf{R_{xx}}\in\mathbb{R}^{NL\times{NL}}$, $\mathbf{R_{ss}}\in\mathbb{R}^{L_{s}\times{L_{s}}}$, and $\mathbf{R_{xs}}\in\mathbb{R}^{NL\times{L_{s}}}$ are the autocorrelation matrix of $\mathbf{X}$, autocorrelation matrix of $\mathbf{S}$, and cross-correlation matrix of $\mathbf{X}$ and $\mathbf{S}$, respectively.
@@ 2.2.1 Non-linear stimulus reconstruction
Another approach is reconstructing the stimulus envelope using a non-linear model (NSR). In this study, we implemented the CNN-based network proposed by Thornton et al.[17], which meets the criteria described above, as a baseline method for the NSR approach. The network was inspired by the EEGNet architecture developed by Lawhern et al.[29], which comprises two main convolutional blocks, one fully connected (FC) classification layer and employs the exponential linear unit (ELU) as a nonlinear activation function, as well as batch normalization (BN) and average pooling. Details of the network can be found in the original study [17]. Another notable study using the NSR approach is by De Taillez et al.[15]. The authors developed a feed-forward neural network comprising a single hidden layer and an output layer with other DL features such as ‘tanh’ activation functions, dropout (DO) [30], and BN. Despite our best efforts to implement and validate this method, its performance was significantly lower compared to the other methods, and therefore, the results are not included here. A similar observation was reported in [24].
@@ 2.2.2 Proposed direct AAD method
Figure 3: The architecture of AADNet. T = length of input data, N = number of input EEG channels, BN = batch normalization, DO = dropout, FC = fully connected, and h = output size of the FC1 layer.
@@ 3.1 Dataset
In this work, we use three datasets recorded using a competing-talker setup, which has been used in previous AAD-related studies to address the AAD problem.
@@ 3.1.2 Dataset II - DTU
3.1.3 Dataset III - KUL
Dataset III [35] was also collected with a speech competing scenario to address the AAD problem in a soundproof, electromagnetically shielded room using a 64-channel BioSemi ActiveTwo system and sampled at a frequency of 8196 Hz. The data comprises 16 normal hearing subjects with 20 trials per subject. The first 8 trials are 6-minute-long, and the remaining 12 trials, which are repetitions of the first trials, are 2-minute-long. During each trial, two different Dutch stories narrated by male speakers were presented simultaneously through a pair of insert phones. The subjects were instructed to attend to the story in one ear while ignoring the one in the other ear. After each trial, subjects were asked a set of muliple-choice questions about the story. There were two stimulus conditions, i.e., ‘dichotic’ and head-related transfer filtered with ±${\pm}90^{\circ}$ along the azimuth direction. The attended ear and stimulus condition were permuted and kept equally distributed over the trials.
Table 2 summarizes the details of the three datasets.
@@ 3.1.3 Dataset III - KUL
DTU 18 60 50 64 Insert earphones, simulated speaker ±${\pm}60^{\circ}$ direction using a head-related transfer function, 2.4 m distance Anechoic, mildly, and highly reverberant
KUL 16 8+4 360/120 64 Insert earphones, simulated speaker ±${\pm}90^{\circ}$ direction using a head-related transfer function and dichotic Electromagnetically shielded room
@@ 3.2 Data pre-processing
In this section, we present the processing steps that were applied on the three presented datasets for the LSR, CCA, and the proposed AADNet models. For the NSR model, we followed the processing steps proposed in the original study [17].
@@ 3.2.1 EEG
The EEG data were first band-pass filtered using a zero-phase FIR filter with cutoffs at 0.5 Hz and 32 Hz to eliminate slow drift and irrelevant high frequencies. This frequency band was also applied in various studies [15, 16, 24] and is reported to be relevant for envelope-based AAD studies. The data for each channel were then downsampled to 64 Hz, re-referenced to the average of all channels (32 channels for the EventAAD dataset and 64 channels for the DTU and KUL dataset), and zero-centered. For the NSR model, the EEG data were band-pass filtered from 0.25 to 36 Hz using a Hamming window, FIR filter, resampled to 125 Hz, and standardized to have zero mean and unit variance. The pre-processing pipeline was implemented using the Python MNE package version 1.2.0 [36] and SciPy package version 1.10.1 [37].